OpenStudy (anonymous):

Given the function f (x) = 7x + 9, find the formula for the inverse, f -1

OpenStudy (johnweldon1993):

for the inverse....just switch the x's and y's you have y = 7x + 9 so x = 7y + 9 and then solve for y again

OpenStudy (anonymous):

so how do i format the formula? I'm clueless as to where to begin.

OpenStudy (johnweldon1993):

dont worry about it :) so you have x = 7y + 9 to find the inverse...you want to isolate y so add 9 to both sides.... you now have?

OpenStudy (johnweldon1993):

subtract 9 from both sides*** sorry about that

OpenStudy (anonymous):

ok... yes. Thank you

OpenStudy (johnweldon1993):

so now you have x - 9 = 7y you want y COMPLETELY by itself...so divide both sides by 7 \[y = \frac{ x-9 }{ 7 }\] this would be your inverse equation :)

OpenStudy (anonymous):

OMG....makes sense...Thank you.

OpenStudy (johnweldon1993):

lol I love it when that happens haha.....you're very welcome :)